OpenStudy (anonymous):

(2X^2-98)/(X^2+4X-21)=

OpenStudy (dumbcow):

\[\frac{2x^{2}-98}{x^{2}+4x-21} = \frac{2(x^{2}-49)}{(x+7)(x-3)} = \frac{2(x+7)(x-7)}{(x+7)(x-3)} = \frac{2(x-7)}{x-3}\]

OpenStudy (dumbcow):

Basically just factor everything and things will cancel

OpenStudy (anonymous):

2 factors out of each term on top. Which gives us a numerator of: 2(x^2-49) Now for the denominator: We know that 7*3 = 21, but we need to find a combination of 7 and 3 that will make 4 when added together. 7+3 = 10 -7 + 3 = -4 -7 - 3 = -10 7 - 3 = 4 So we are looking for x + 7 and x - 3 So our denominator factors to (x+7)(x-3) We have the new equation of 2(x^2-49)/(x+7)(x-3) Now we are looking for a factored equation which will become x^2-49. We know that 7 * 7 is 49. If we add 7 and -7 we have 0, which is the inner term above. (x^2+0x-49) So we now have: 2(x+7)(x-7)/(x+7)(x-3) Now the x+7 above and below will cancel. 2(x-7)/(x-3) Hope this helps. http://www.tutorsean.net